The radius of a sphere is increased by 2.5%.find the percentage increase in the volume
Answer by Alan3354(69443)

The radius of a sphere is increased by 2.5%.find the percentage increase in the volume
-----------
= Vol*1.025^3 = 1.076890625*Vol
--> ~ 7.689% increase